# Information-Driven Design of Imaging Systems

### v4

@@ Abstract
In modern imaging systems that computationally process raw measurements before or instead of human viewing, information content matters more than visual appearance. However, developing information estimators that can handle the complexity of real-world measurements yet remain practical enough for widespread use has proven challenging. We introduce a data-driven approach for estimating mutual information between unknown objects and their noisy measurements. Our technique fits probabilistic models to measurements and their noise processes, quantifying information content without requiring ground truth data or making assumptions about object structure. We validate our approach across diverse applications—color photography, radio astronomy, lensless imaging, and microscopy—demonstrating that information estimates reliably predict system performance. Finally, we introduce Information-Driven Encoder Analysis Learning (IDEAL), which optimizes imaging systems to maximize information capture. Our work unlocks information theory as a powerful, practical tool for analyzing and designing imaging systems across a broad range of applications.
A video summarizing this work can be found at:
@@ 1 Introduction
Imaging systems have traditionally been designed to mimic the human eye and produce visually interpretable pictures. Today, however, algorithms process most imaging data before or instead of human viewing.
Unlike human vision, which is inherently sensitive to specific visual patterns [1], algorithms like modern deep neural networks can extract meaning from information encoded in many forms. Computational imaging systems [2] often leverage these capabilities by extracting image features that are not visually apparent in raw measurements.
@@
This shift fundamentally changes what limits performance: not how well systems encode information into human-perceptible patterns, but how much information they encode, regardless of its form. Quantifying information content thus offers a powerful tool for designing better ways to capture data.
The information content of measurements is often assessed indirectly after computational decoding, which presents significant limitations. In this approach, a decoding algorithm processes raw measurements to perform downstream tasks (e.g., reconstructing the original object), the result of which is compared to ground truth data to calculate performance metrics [2, 3, 4]. Since this approach evaluates the encoder (hardware) and decoder (algorithm) simultaneously, it leaves ambiguity about whether to focus improvement efforts on hardware or algorithms. For example, many decoders use assumptions about the original object that affect performance yet have no relation to measurement quality. Additionally, this method requires ground truth data for comparison, which usually cannot be obtained in experimental settings.
Alternatively, the information content in raw measurements can be directly quantified: mutual information quantifies how much object information survives the encoding and measurement process [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 3]. This approach is independent of measurement visual appearance or downstream processing. It unifies traditionally separate metrics like resolution and signal-to-noise ratio, and accounts for effects such as quantization and sampling [9, 17, 30]. As a result, mutual information enables objective evaluation of design tradeoffs affecting multiple aspects of system performance (Sec. S1.2).
Despite these benefits, accurately estimating information content in real imaging systems has proven challenging. Previous approaches fall into two categories. The first calculates theoretical limits [5, 7, 8, 9, 10, 15, 20, 29] that, while powerful for communication systems (in which the encoder can map any input to any output), produce overly optimistic bounds that fail to account for the physical constraints of real imaging systems (Fig. S6). The second develops analytical probabilistic models of objects that work in specific cases but require domain-specific assumptions that are difficult to validate experimentally.
In this work, we introduce an information estimation framework that achieves the best of both worlds: it inherently captures physical constraints by evaluating real measurements, while avoiding domain-specific model development. Our approach requires just two inputs: a dataset of measurements and a characterization of system noise. First, we estimate the measurements’ entropy by fitting a probabilistic model to them. Then we subtract the estimated entropy due to noise, giving a measure of the information about the object that is passed into the measurement. By eliminating the need for assumptions about the imaged object or imaging system, our approach can evaluate diverse imaging modalities with complicated performance tradeoffs under real-world conditions.
In the following sections, we describe our information estimation framework and experimentally validate its performance. We then demonstrate its practical utility across diverse imaging domains: color photography, black hole radio astronomy, non-traditional optics in lensless cameras, and biological microscopy. Finally, we introduce Information-Driven Encoder Analysis Learning (IDEAL), which transforms our evaluation framework into a design tool by combining it with gradient-based optimization. This enables automated design of imaging systems optimized for maximum information capture.
2 Formulation of information estimator
Mutual information estimation requires modeling objects, images, and measurements as probability distributions (Fig. 1). For example, a probability distribution over tree objects captures the relative likelihood of each possible variation in size, shape, and species. When we apply an encoder (e.g. capturing an image using a camera with a particular lens and position) to each possible object, we obtain a noiseless image distribution, such as the images of each possible tree on a 2D sensor. Random noise then corrupts these noiseless images to create the measurement distribution we actually observe [5, 14, 22, 31, 32] (Sec. S1).
Within this probabilistic framework, we aim to quantify how effectively different encoders preserve information about the object. Each encoder represents a different system design; “good” encoders will preserve information throughout the system. Information can be lost both during the encoding process (e.g. due to finite resolution or sampling) or during the measurement process (e.g. due to noise). The mutual information between object and measurement distributions quantifies how much object information survives this entire process.
@@ 2 Formulation of information estimator
Figure 1: Information estimation for imaging systems. An encoder (i.e., an optical system) maps unknown objects to noiseless images $\mathbf{X}$, which noise corrupts to produce measurements $\mathbf{Y}$. Our information estimator takes these measurements and a noise model as inputs to quantify how well different objects can be distinguished. We can use the estimate $\hat{I}(\mathbf{X};\mathbf{Y})$ to evaluate existing systems or guide the design of new ones.
@@
Mathematically, we represent an image with $D$ pixels as a vector $\mathbf{x}=(x_{1},x_{2},\dots,x_{D})$, where $x_{k}$ is the energy at the $k$th pixel of an image sensor. The probability distribution over images is given by the $D$-dimensional joint distribution over pixels $p(\mathbf{x})=p(x_{1},x_{2},\dots,x_{D})$, which gives the probability of each possible image. Similarly, we represent the noisy measurement as $\mathbf{y}=(y_{1},y_{2},\dots,y_{D})$ with probability density $p(\mathbf{y})$.
Our goal is to estimate the mutual information between object and measurement, $I(\mathbf{O};\mathbf{Y})$, which quantifies how well measurements can distinguish different objects (Sec. S2.4). To accomplish this in experimental settings, we must work with only a dataset of measurement samples $\{\mathbf{y}^{\left(i\right)}\}_{1}^{N}$, without making any assumptions about the underlying object or image distributions.
We cannot directly estimate $I(\mathbf{O};\mathbf{Y})$ without access to the object distribution. Fortunately, for most imaging systems, we can instead focus on $I(\mathbf{X};\mathbf{Y})$ (the mutual information between noiseless image and noisy measurement distributions). This approach simplifies our task because it depends only on the noiseless image and noisy measurement distributions. This simplification is valid for imaging systems that always produce the same noiseless image for a given object. In such cases, all information in a noiseless image comes from the object, and any information in an image that is not destroyed by noise during the measurement process must have also come from the object. Mathematically, $I(\mathbf{O};\mathbf{Y})=I(\mathbf{X};\mathbf{Y})$. (In situations where $I(\mathbf{O};\mathbf{Y})\neq I(\mathbf{X};\mathbf{Y})$ because of an additional source of randomness, such as a changing encoder, the estimator can be adapted to account for this using additional system calibration data (Sec. S1.5.1)).
@@
Estimating mutual information between high-dimensional variables from samples is notoriously challenging [33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Without strong assumptions, estimation requires sample sizes that grow exponentially with dimensionality, suffers from high bias and variance, and lacks formal guarantees  [43, 38]. While numerous methods exist for estimating using paired samples $\{\mathbf{x}^{\left(i\right)},\mathbf{y}^{\left(i\right)}\}_{1}^{N}$, for physical instruments like imaging systems, we face a fundamental limitation: we typically only have access to noisy measurements $\{\mathbf{y}^{\left(i\right)}\}_{1}^{N}$.
@@
However, imaging systems offer a key simplification: the noise process $p(\mathbf{y}\mid\mathbf{x})$ [32] is often well-characterized through physical models or empirical calibration. This enables decomposing $I(\mathbf{X};\mathbf{Y})$ into entropy terms that can be estimated from noisy measurements alone:
@@
These entropies quantify the amount of variation, or diversity, in each distribution; a measure of how unpredictable the pixel values are. Here, $H(\mathbf{Y})$ quantifies the total entropy in the measurements, arising from both object information and noise. To isolate object information, we subtract $H(\mathbf{Y}|\mathbf{X})$, the entropy due to noise alone. This difference, $I(\mathbf{X};\mathbf{Y})$, thus gives our desired metric – the measurement entropy that comes from object variations, which quantifies how well the encoder preserves object information.
The two terms in Eq. 1 can be written explicitly in terms of probability distributions:
@@
The second term (Eq. 3) can be simplified when the noise at each pixel is independent of other pixels conditioned on the noiseless image—a common assumption in models of imaging systems [32] (Sec. S2.3). This reduces the computation of $H(\mathbf{Y}|\mathbf{X})$ to a sum $\sum_{i=1}^{D}H(\mathrm{Y}_{i}|\mathrm{X}_{i})$ of the conditional entropies at each pixel (Sec. S2.3.1), avoiding an intractable high-dimensional integral.
@@
The first term (Eq. 2) presents a greater challenge since we don’t know the true distribution of measurements $p(\mathbf{y})$. To address this, we fit a parametric model $p_{\theta}(\mathbf{y})$ to a training set of $N$ measurements $\{\mathbf{y}^{\left(i\right)}\}_{i}^{N}$, then estimate $H(\mathbf{Y})$ using the cross-entropy upper bound on a held-out test set of $M$ additional measurements:
$H(\mathbf{Y})\leq\hat{H}(\mathbf{Y})$ $=\mathbb{E}\left[-\log p_{\theta}(\mathbf{Y})\right]$ (4)
$\approx-\frac{1}{M}\sum_{i=1}^{M}\log p_{\theta}(\mathbf{y}^{\left(i\right)})$ (5)
This approach has been shown to reliably estimate entropy in complex real-world distributions, even without direct knowledge of the underlying generative process [44].
@@
Furthermore, since this cross-entropy (between the true and estimated distributions) is always larger than the true entropy, this provides a principled way to compare different models: those that achieve lower cross-entropy values produce more accurate estimates. Subtracting our estimate of $H(\mathbf{Y}|\mathbf{X})$ from this upper bound on $H(\mathbf{Y})$ yields our estimate of the mutual information $\hat{I}(\mathbf{X};\mathbf{Y})$.
3 Implementation and validation
Practically implementing our theoretical framework for information estimation requires choosing an appropriate probabilistic model to fit to the measurement distributions. This choice carries an inherent trade-off between model expressivity and computational efficiency. We explored three model architectures that offer different compromises for this trade-off.
Our models differ in two key assumptions: (1) whether the measurement distribution is constrained to be multivariate Gaussian, which simplifies computation but limits expressivity for non-Gaussian statistics, and (2) whether the model is translation-invariant (stationary), meaning statistical relationships between pixels depend only on their relative positions rather than absolute locations. Stationarity significantly reduces parameter count but cannot capture position-dependent artifacts like fixed sensor defects or illumination gradients.
The three models, in order from least to most expressive, are:
  • A stationary Gaussian process, which assumes both a Gaussian distribution and translation invariance. It has the fewest parameters (Fig. S10), making it useful with limited data, particularly when the underlying statistics are in fact stationary. This model can be fit in 0.1 seconds using $20\times 20$ image patches on an NVIDIA RTX A6000 GPU. We developed an optimization procedure to improve its numerical stability and data efficiency (Fig. S11) at the cost of longer ($\sim$ 10s) training times.
  • A Gaussian process with no constraints on its covariance matrix (i.e. it is non-stationary). This provides greater flexibility while maintaining good computational efficiency ($\sim$ 0.1s to fit), but still lacks the ability to accurately fit many real-world distributions.
  • An autoregressive PixelCNN [45, 46, 47] model, which can capture complex non-Gaussian measurement distributions, but requires the most data and computation time ($\sim$ 100s) (Fig. S17).
Which model to choose depends on the specific application: the full Gaussian model is fast for large datasets, PixelCNN is the most accurate when data are abundant, and the stationary Gaussian model can be used when data are scarce. When computational resources allow, the best approach is to train multiple models and select the lowest entropy estimate, since tighter entropy bounds yield better information estimates.
To increase the practicality of fitting models, we processed images as smaller patches of full images. Though larger patches capture more of the statistical relationships between distant pixels in images and can yield more accurate estimates (Fig. S12), smaller patches increase the number of samples for training while also reducing model complexity and computational burden.
We validated the expressiveness and accuracy of our models through a series of experiments using both simulated and real-world data. First, we used simulated measurement data from an analytic model with a known ground truth mutual information: a stationary Gaussian process with additive noise. All three models converged to the correct value if given sufficient data (Fig. S16a). For real-world non-Gaussian data, PixelCNN achieved tighter bounds than Gaussian models (Fig. S16b,c) as demonstrated by its ability to better capture the measurements’ spatial structure  (Fig. S14). This advantage is particularly clear on the MNIST dataset of handwritten digits, where pixels tend to be either black or white, creating a bimodal distribution that Gaussian models cannot model well (Fig. S15). Applied to real microscopy data at different photon counts (Fig. S13), our estimators captured the expected increase in information content with improved signal-to-noise ratio across different contrast modalities. Further validation on optical imaging systems limited by photon shot noise (i.e. Poisson-distributed noise) demonstrated that the conditional entropy $H(\mathbf{Y}|\mathbf{X})$ can be estimated accurately using only noisy measurements, except at very low photon counts (Fig. S18).
4 Information estimation on imaging applications
Though theoretical results show that mutual information sets a limit on achievable performance for many types of decoders [48, 49, 50, 51], real systems deviate from theoretical assumptions and information estimates are imperfect. To further validate our information estimator, we tested its ability to predict the performance of four real imaging systems. These imaging systems use a range of post-processing decoders to obtain the final image, from traditional signal processing approaches to modern neural networks. Typically, system evaluation computes error relative to ground truth data after applying the decoder algorithm to perform tasks like image reconstruction or classification. Our estimator does not require ground truth data and can still predict the performance of the imaging system for various different objects, noise regimes, tasks, and performance metrics, independent of which type of decoder is used. This validates information estimation as a reliable way to evaluate performance without the need for application-specific decoders or ground truth data.
Color photography
Our first application example involves analyzing color filter arrays for digital photography. Digital cameras capture color images by placing a color filter in front of each pixel, such that it receives only red, green, or blue light; this is the encoding process. Then, post-processing algorithms interpolate this data and reconstruct a full color image with all three colors at each pixel; this is the decoding process, referred to as “demosaicing.” The widely used Bayer pattern arranges filters in a repeating 2$\times$2 grid with two green pixels, one red, and one blue (Fig. 2a, bottom). Recent work has suggested that learned designs (encoders) which include “white” pixels (no color filter) can yield higher fidelity reconstructions when combined with modern demosaicing algorithms [52, 53].

### v5

@@ Abstract
Imaging systems have traditionally been designed to mimic the human eye and produce visually interpretable measurements. Modern imaging systems, however, process raw measurements computationally before or instead of human viewing. As a result, the information content of raw measurements matters more than their visual interpretability. Despite the importance of measurement information content, current approaches for evaluating imaging system performance do not quantify it: they instead either use alternative metrics that assess specific aspects of measurement quality or assess measurements indirectly with performance on secondary tasks.
We developed the theoretical foundations and a practical method to directly quantify mutual information between noisy measurements and unknown objects. By fitting probabilistic models to measurements and their noise characteristics, our method estimates information by upper bounding its true value. By applying gradient-based optimization to these estimates, we also developed a technique for designing imaging systems called Information-Driven Encoder Analysis Learning (IDEAL). Our information estimates accurately captured system performance differences across four imaging domains (color photography, radio astronomy, lensless imaging, and microscopy). Systems designed with IDEAL matched the performance of those designed with end-to-end optimization, the prevailing approach that jointly optimizes hardware and image processing algorithms. These results establish mutual information as a universal performance metric for imaging systems that enables both computationally efficient design optimization and evaluation in real-world conditions.
A video summary of this work can be found at:
@@ 1 Introduction
The increasing utilization of computational processing in imaging systems removes the constraint that they must produce human-interpretable measurements. Traditional imaging system designs produce measurements that contain the specific visual patterns recognizable to human observers [fassHumanSensitivityMutual]. However, these patterns are unnecessary for modern algorithms like deep neural networks, which can utilize information in measurements regardless of how it appears: they can either transform measurements into forms that make sense to humans, as computational imaging systems do [bhandariComputationalImaging2022], or they can analyze measurements directly and bypass humans altogether.
@@
This shift fundamentally changes what limits performance: not how well systems encode information into human-perceptible patterns, but how much information they encode, regardless of its visual appearance.
Existing metrics to quantify hardware performance fail to capture information content comprehensively. Traditional measurement metrics like resolution, signal-to-noise ratio, sampling rate, and field-of-view characterize individual system aspects separately, making it difficult to assess their combined effect on overall performance or compare systems that make different trade-offs among these factors.
As a result, performance is often assessed by computationally decoding measurements to perform secondary tasks. This decoder-based approach uses tasks like image reconstruction or object classification. It quantifies success by comparing algorithm outputs against ground truth using metrics like mean squared error or classification accuracy [bhandariComputationalImaging2022, yuanGeometricDeepOptical2023, wetzsteinInferenceArtificialIntelligence2020]. This approach furthermore enables system design in simulation through “end-to-end” methods that jointly optimize both encoder parameters (like lens shapes or sensor configurations) and decoder algorithms (like image reconstruction). [sitzmannEndtoendOptimizationOptics2018, sunLearningRank1Diffractive2020, metzlerDeepOpticsSingleShot2020, sunEndtoendComplexLens2021, tsengNeuralNanoopticsHighquality2021, tsengDifferentiableCompoundOptics2021, debFourierNetsEnableDesign2022].
However, decoder-based evaluation and end-to-end optimization have a few limitations. First, decoder-based evaluation metrics require full ground truth knowledge of the object or some of its properties, which limits their application to simulations and laboratory settings that may not generalize to real-world conditions. Second, this joint evaluation conflates encoder and decoder performance, since low performance could result from either insufficient encoded information or the decoder’s failure to utilize it. Finally, neural network-based decoders require substantial compute and may create backpropagation challenges in end-to-end optimization [debFourierNetsEnableDesign2022, yangImageQualityNot2023a].
Directly quantifying information in measurements provides an alternative approach that avoids the limitations of decoders. Mutual information quantifies how much object information survives the encoding and measurement processes [fellgettAssessmentOpticalImages1955, toraldodifranciaResolvingPowerInformation1955, difranciagtDegreesFreedomImage1969, bershadResolutionOpticalChannelCapacity1969, falesImagingSystemDesign1984, coxInformationCapacityResolution1986, linskerSelforganizationPerceptualNetwork1988, linskerApplicationPrincipleMaximum, vandyckUltimateResolutionInformation1992, huckInformationTheoryVisual1997, neifeldInformationResolutionSpace1998, osullivanInformationtheoreticImageFormation1998, gartenbergInformationMetricDesign2000, prasadInformationCapacitySeeinglimited2000, caoInformationTheoreticAssessment2001, wagnerInformationTheoreticalOptimization2003, neifeldInformationoptimizedImagingSystem2001, ashokInformationbasedAnalysisSimple2003, neifeldTaskspecificInformationImaging2007, wuOptimalPhaseTransitions2012, ashokInformationOptimalCompressive2013, huangInformationOptimalCompressive2015, gureyevSignaltonoiseSpatialResolution2018, narimanovResolutionLimitLabelfree2019, willomitzerSingleShot3DSensing2019, yuanGeometricDeepOptical2023]. Equivalently, it quantifies how much a measurement reduces an observer’s uncertainty about the object.
Mutual information also unifies traditionally separate aspects of measurement quality, enabling meaningful comparison between systems that make different trade-offs among these factors (Sec. S1.2). For example, it captures the combined effect of resolution, signal-to-noise ratio, sampling, and quantization on information preservation [falesImagingSystemDesign1984, gartenbergInformationMetricDesign2000, ashokInformationOptimalCompressive2014].
Despite these benefits, practical estimation of mutual information has been challenging. Previous approaches fall into two categories, each with significant limitations.
The first approach overestimates information because it models an imaging system as an unconstrained communication system [fellgettAssessmentOpticalImages1955, difranciagtDegreesFreedomImage1969, bershadResolutionOpticalChannelCapacity1969, falesImagingSystemDesign1984, coxInformationCapacityResolution1986, neifeldInformationResolutionSpace1998, wagnerInformationTheoreticalOptimization2003, willomitzerSingleShot3DSensing2019]. While communication systems assume encoders can transform any input into any output without physical constraints, imaging systems cannot achieve this because lenses, sensors, and other hardware impose fundamental limitations. These theoretical calculations can therefore produce highly inaccurate estimates (Fig. S6).
The second approach explicitly models the objects being imaged, relying on assumptions about object properties that are difficult to verify in practice [ashokInformationbasedAnalysisSimple2003, neifeldTaskspecificInformationImaging2007, wuOptimalPhaseTransitions2012, ashokInformationOptimalCompressive2013, huangInformationOptimalCompressive2015, gureyevSignaltonoiseSpatialResolution2018]. These object models must also be developed case-by-case for each object type, limiting their generality.
To address these limitations, we developed a framework for estimating mutual information directly from measurement of unknown objects. Our approach requires only two inputs: a dataset of measurements and a model of the system’s noise. By fitting models to actual measurements rather than assuming object properties or ignoring encoders’ physical constraints, the method readily adapts to diverse imaging modalities and performance trade-offs. Our framework enables computationally efficient optimization when designing systems through simulation, as well as quantitative evaluation of real imaging systems operating in the field.
This paper is organized as follows: Section 2 presents the theoretical framework for estimating mutual information from measurements. Section 3 describes implementation through three probabilistic models and validates their accuracy. Section 4 demonstrates that information estimates predict system performance across four imaging applications: color photography, radio astronomy, lensless imaging, and microscopy. Section 5 introduces our method for automated imaging system design through gradient-based optimization, Information-Driven Encoder Analysis Learning (IDEAL). Section 6 considers implications and future research directions.
2 Mathematical framework
This section shows how to overcome the inherent statistical and computational challenges of mutual information estimation by progressively decomposing the estimation problem using key properties of imaging systems. First, deterministic encoding makes it possible to focus on the relationship between clean images and noisy measurements rather than the more complex object-measurement relationship. Second, the known statistical structure of noise in imaging systems enables decomposing the estimation problem into two simpler subproblems. Finally, the constraint of only having measurement data (on real imaging systems) informs the selection of estimation approaches for each subproblem.
To formalize this approach, we model imaging systems as a chain of probability distributions (Fig. 1). Consider photographing trees: a probability distribution over objects captures how trees vary in size, shape, and species. An encoder (a camera with particular lens and position) maps each possible tree to a noiseless image on the sensor. Detection noise corrupts these ideal images, creating the measurements that sensors actually record [fellgettAssessmentOpticalImages1955, huckInformationTheoryVisual1997, ashokInformationbasedAnalysisSimple2003, changInformativeSensing2009, goodmanStatisticalOptics2015] (Sec. S1).
Mutual information provides a metric for comparing encoder designs by quantifying how well measurements can distinguish different objects. Information losses occur at two stages: during encoding (e.g. from finite resolution or sampling) and during measurement (from noise).
@@ 2 Mathematical framework
Figure 1: Information estimation for imaging systems. An encoder (e.g., an optical system) maps unknown objects to noiseless images $\mathbf{X}$. Noise corrupts these images to produce measurements $\mathbf{Y}$. The information estimator uses these measurements and a noise model to quantify how well measurements distinguish noiseless images (and thus objects under deterministic encoding). The estimate $\hat{I}(\mathbf{X};\mathbf{Y})$ enables both encoder evaluation and design optimization.
@@
Mathematically, we model images and measurements as probability distributions over vectors. An image with $D$ pixels is a vector $\mathbf{x}=(x_{1},x_{2},\dots,x_{D})$, where $x_{k}$ represents the energy at the $k$th pixel. The probability distribution $p(\mathbf{x})=p(x_{1},x_{2},\dots,x_{D})$ describes the likelihood of each possible image. Similarly, a noisy measurement is a vector $\mathbf{y}=(y_{1},y_{2},\dots,y_{D})$ with distribution $p(\mathbf{y})$.
Quantifying object information in measurements becomes tractable through the widely-used assumption of deterministic imaging systems. The object-measurement mutual information $I(\mathbf{O};\mathbf{Y})$ requires working with the joint distribution of objects and measurements, which is unknown in real systems. However, deterministic encoders, which always produce the same noiseless image for each object, enable estimation of $I(\mathbf{X};\mathbf{Y})$ instead. This deterministic encoder assumption applies broadly in practice and holds by construction in simulated design.
Intuitively, this equality holds because all variations in noiseless images come from variations in objects when encoders are deterministic. Thus, any image information that survives the measurement process must have also come from the object. Mathematically:
Deterministic Information Transfer Theorem. For a Markov chain $\mathbf{O}\to\mathbf{X}\to\mathbf{Y}$, if $\mathbf{X}=f(\mathbf{O})$ for some deterministic function $f$, then $I(\mathbf{O};\mathbf{Y})=I(\mathbf{X};\mathbf{Y})$.
(See supplement for proof (Sec. S1.1)).
@@
Even the more accessible $I(\mathbf{X};\mathbf{Y})$ estimation problem requires handling the high-dimensional joint distribution $p(\mathbf{x},\mathbf{y})$. Estimating mutual information between high-dimensional variables from samples is notoriously challenging [kraskovEstimatingMutualInformation2004a, darbellayEstimationInformationAdaptive1999, rossMutualInformationDiscrete2014, pooleVariationalBoundsMutual2019, goldfeld2021sliced, Song2020Understanding, czyz2023beyond, gowriMIEstimateLearnedReps2024, letiziaMIEstimate2024, butakovMI_EstNormFlow2024]. Without strong assumptions, estimation requires sample sizes that grow exponentially with dimensionality, suffers from high bias and variance, and lacks formal guarantees  [mcallesterFormalLimitationsMeasurement2018, Song2020Understanding].
Furthermore, the majority of existing estimation methods assume access to a dataset of paired samples $\{\mathbf{x}^{\left(i\right)},\mathbf{y}^{\left(i\right)}\}_{1}^{N}$. While these are available in simulated systems, real systems only provide noisy measurements $\{\mathbf{y}^{\left(i\right)}\}_{1}^{N}$ without corresponding noiseless images.
@@
Fortunately, imaging systems have well-characterized noise processes $p(\mathbf{y}\mid\mathbf{x})$ from physical models or calibration [goodmanStatisticalOptics2015]. This knowledge, combined with the deterministic encoder property, enables decomposing $I(\mathbf{X};\mathbf{Y})$ into entropy terms that can be estimated from noisy measurements alone:
@@
$H(\mathbf{Y})$ quantifies total variation in measurements from both objects and noise. $H(\mathbf{Y}|\mathbf{X})$ quantifies variation from noise alone. Their difference $I(\mathbf{X};\mathbf{Y})$ isolates noiseless image-induced variation in measurements, which equals the object information preserved when encoders are deterministic.
These entropy terms can be expressed as expectations:
@@
The $H(\mathbf{Y}|\mathbf{X})$ term further simplifies under the assumption of independent noise across pixels, reducing the conditional entropy $H(\mathbf{Y}|\mathbf{X})$ from a high-dimensional integral to a sum of single-pixel entropies: $H(\mathbf{Y}|\mathbf{X})=\sum_{i=1}^{D}H(\mathrm{Y}_{i}|\mathrm{X}_{i})$ (Sec. S2.3.1). Imaging systems commonly satisfy this assumption [goodmanStatisticalOptics2015].
@@
The $H(\mathbf{Y})$ term requires learning from data. The true distribution $p(\mathbf{y})$ is unknown but its entropy $H(\mathbf{Y})$ can nonetheless be upper bounded by fitting a parametric model $p_{\theta}(\mathbf{y})$ to a training set of $N$ measurements $\{\mathbf{y}^{\left(i\right)}\}_{1}^{N}$ and computing the cross-entropy on a held-out test set of $M$ additional measurements:
$H(\mathbf{Y})\leq\mathbb{E}\left[-\log p_{\theta}(\mathbf{Y})\right]\approx-\frac{1}{M}\sum_{i=1}^{M}\log p_{\theta}(\mathbf{y}^{\left(i\right)})$ (4)
This upper-bounding approach reliably estimates entropy in complex real-world distributions without knowing the true probability [henighanScalingLawsAutoregressive2020].
@@
Furthermore, cross-entropy provides a principled way to compare models. Since cross-entropy between true and estimated distributions always exceeds true entropy, models achieving lower cross-entropy values produce more accurate estimates.
Subtracting our estimate of $H(\mathbf{Y}|\mathbf{X})$ from this upper bound on $H(\mathbf{Y})$ yields our estimate of the mutual information $\hat{I}(\mathbf{X};\mathbf{Y})$.
3 Probabilistic models for information estimation
Practical implementation of our framework requires choosing probabilistic models to fit both measurements and noise processes. For measurements, users face a fundamental trade-off: expressive models capture complex distributions accurately but demand substantial data and compute, while simpler models train faster using fewer resources but may overestimate information. To address this trade-off, we developed and tested three models spanning the expressivity-efficiency spectrum: a stationary Gaussian process, a full Gaussian process, and an autoregressive PixelCNN model [oordPixelRecurrentNeural2016, vandenoordConditionalImageGeneration2016, salimansPixelCNNImproving2017]. For noise modeling, users must select an appropriate model and (on real systems) fit its parameters using only noisy measurements.
This section details the three measurement models and validates their performance through experiments that establish practical guidance for model selection in different scenarios.
Measurement model assumptions. Our models differ in two key assumptions: whether measurements follow a multivariate Gaussian distribution and whether pixel statistics are translation-invariant (stationary). Violated assumptions yield worse fits that overestimate information. For example, Gaussian models cannot capture many types of statistical structure that are present in real measurement distributions, such as the bimodal statistics of measurements of MNIST handwritten digits (Fig. S15). Meanwhile, stationary models fail with position-dependent artifacts like sensor defects or illumination gradients.
We processed images as patches rather than full images for computational efficiency. Larger patches capture more spatial relationships (Fig. S12) but require more computation; smaller patches provide more training samples and permit smaller models.
Model types. We tested three models with different expressivity-efficiency trade-offs.
Table 1: Comparison of probabilistic models for information estimation. Training times measured on $20\times 20$ patches using an NVIDIA RTX A6000 GPU. Stationary Gaussian training time range reflects optional optimization for numerical stability.
Model Data Efficiency Training Time Expressivity Gaussian Stationary
Stationary Gaussian Highest $\sim$0.1–10s Lowest
Gaussian (full cov.) Medium $\sim$0.1s Medium
PixelCNN Lowest $\sim$100s Highest
Estimator consistency. We validated convergence behavior on both simulated data with known ground truth and real imaging data. On simulated data from a stationary Gaussian process with additive noise, all three models converged to the analytically-calculated mutual information, with the stationary Gaussian converging fastest since the test data matched its assumptions (Fig. S16a). On real imaging data, all models converged to stable, though different, values. PixelCNN consistently achieved the lowest entropy bounds, indicating the most accurate estimates, while Gaussian models plateaued at higher values due to their inability to capture non-Gaussian statistics (Fig. S16b,c). Based on these convergence results, we used $\sim$1,000–10,000 patches of $\sim 20\times 20$ pixels for subsequent experiments, balancing computational efficiency with estimation accuracy.
Model selection. Based on our validation results, we recommend selecting models according to available resources and data characteristics. The stationary Gaussian provides reliable estimates with limited data, PixelCNN maximizes accuracy when computational resources and data are abundant, and the standard Gaussian balances ease of computation with reliable performance. While PixelCNN generally provides the most accurate estimates, the performance difference may be smaller for data with approximately Gaussian statistics, making simpler Gaussian models reasonable when computational resources are limited. Fitting multiple models and selecting the lowest entropy estimate yields the most reliable results when feasible.
Noise model validation. Noise processes in imaging systems proved far simpler to model than measurement distributions, making $H(\mathbf{Y}|\mathbf{X})$ estimation straightforward. While this required identifying the correct noise model and fitting its parameters, the structure of imaging noise allows the framework to remain robust to somewhat inaccurate noise modeling. Object-independent noise (thermal, dark current) affects all measurements equally, so any model error simply shifts all estimates by the same constant, preserving encoder rankings. Object-dependent noise (shot noise) arises from well-understood quantum processes with established mathematical models, facilitating accurate characterization.
We tested one key challenge specific to shot noise and confirmed our framework handles it successfully: shot noise parameters depend on the noiseless image that real systems cannot observe. We validated that accurate entropy estimation remained possible using only noisy measurements. Our estimators succeeded across a wide range of photon counts, deviating from the true value only below $\sim$20 photons per pixel (Fig. S18).
4 Validation across imaging applications
Theoretical results establish that mutual information limits achievable decoder performance [clarksonShannonInformationReceiver2016, clarksonShannonInformationROC2015a, guoMutualInformationMinimum2005a, guoMutualInformationConditional2008], but estimates of information inevitably contain error. If information estimates accurately capture what limits real decoder performance, they should correlate with decoder-based evaluation across diverse imaging systems. We tested this relationship by comparing information estimates against decoder performance across four imaging systems spanning different modalities, noise regimes, and tasks, with decoders ranging from classical signal processing to modern neural networks.
Across all four systems, information estimates predicted decoder performance, validating their use for additional settings where decoders face computational challenges or cannot be used because ground truth data is unavailable.
Color photography.
Digital cameras encode color on monochrome sensors using color filter arrays in front of their pixels. These arrays filter incoming light so that each pixel (traditionally) detects only red, green, or blue light, and demosaicing algorithms decode the raw measurements to reconstruct full-color images at the full sensor resolution. Beyond the traditional Bayer pattern (a $\times2$ grid with two green, one red, and one blue pixel) (Fig. 2a, bottom), recent end-to-end learned designs incorporate white (i.e. no color filter) pixels together with neural-network-based decoders [henzDeepJointDesign2018, chakrabartiLearningSensorMultiplexing] to improve reconstruction quality.